The speed of a car is 10 miles per hour more than the speed of a truck. The distance traveled by the car in 5 hours is the same as the distance traveled by the truck in 6 hours. Find the speed of each vehicle.
Answer by lwsshak3(11628)

The speed of a car is 10 miles per hour more than the speed of a truck. The distance traveled by the car in 5 hours is the same as the distance traveled by the truck in 6 hours. Find the speed of each vehicle.
let x=speed of truck
x+10=speed of car
distance =travel time*speed
6x=5(x+10)
6x=5x+50
x=50
x+10=60
speed of truck=50 mph
speed of car=60 mph